OpenStudy (anonymous):

Choose the correct simplification of the expression (−2x + 10y)(4x − 6y). −8x2 − 52xy + 60y2 −8x2 + 52xy − 60y2 8x2 + 52xy − 60y2 −8x2 + 28xy − 60y2

jhonyy9 (jhonyy9):

multiplie term by term -2x*4x= -2x*(-6y)= 10y*4x= 10y*(-6y)=